OpenStudy (anonymous):

Let y =(root5(x^8+5)). Find values A and B so that (dy)/(dx) = A(x^8+5)^B(C x^D)

myininaya (myininaya):

\[y=\sqrt[5]{x^8+5}\] \[y=(x^8+5)^\frac{1}{5}\] Use the chain rule to find y'

OpenStudy (anonymous):

what if I need to find A,B,C, and D? what do I plug in?

myininaya (myininaya):

You need to find y' first to find A,B,C,and D

OpenStudy (anonymous):

I havent learned the chain rule yet

myininaya (myininaya):

\[[f(g(x))]'=f'(g(x)) \cdot g'(x) \text{ chain rule }\]

myininaya (myininaya):

Well you need the chain rule to do this problem.

myininaya (myininaya):

what is g(x) and what is f(x) if \[f(g(x))=(x^8+5)^\frac{1}{5}\]

OpenStudy (anonymous):

g(x)= 8x f(x)= 5^1/5

myininaya (myininaya):

No g(x) is the inside function and f(x) is the outside function

OpenStudy (anonymous):

oh ok g(x)= 8x^7

myininaya (myininaya):

\[g(x)=x^8+5 ; f(x)=x^\frac{1}{5}\] \[g'(x)=8x^7+0 ; f(x)=\frac{1}{5}x^{\frac{1}{5}-1}=\frac{1}{5}x^\frac{-4}{5}\]

OpenStudy (anonymous):

so I plug this in in the equation now

myininaya (myininaya):

this one \[y'=[f(g(x))]'=f'(g(x)) \cdot g'(x) \text{ chain rule } \] yes

OpenStudy (anonymous):

1/5x^-4/5

myininaya (myininaya):

\[y'=f'(x^8+5) \cdot 8x^7\]

OpenStudy (anonymous):

now just multiply, right?

myininaya (myininaya):

you need to plug x^8+5 into f'

OpenStudy (anonymous):

oh ok

myininaya (myininaya):

\[g'(x)=8x^7+0 ; f'(x)=\frac{1}{5}x^{\frac{1}{5}-1}=\frac{1}{5}x^\frac{-4}{5}\]

myininaya (myininaya):

f' is that second one I wrote down just now

myininaya (myininaya):

\[y'=f'(x^8+5) \cdot 8x^7=\frac{1}{5}(x^8+5)^\frac{-4}{5} \cdot 8x^7\]

myininaya (myininaya):

now you should be able to put it in the form above to get your A,B,C, and D

OpenStudy (anonymous):

i'll try it

myininaya (myininaya):

actually it is already in that form

OpenStudy (anonymous):

so just look for the A,B,C,D then right?

myininaya (myininaya):

yes